OpenStudy (anonymous):

What is the discriminant for the quadratic expression? 18x2 + 11x A. 49 B. 121 C. 193 D. 280

OpenStudy (anonymous):

discriminant = b² - 4ac y = ax² + bx + c is the quadratic equation, so you have a = 18, b = 11, c = 0

OpenStudy (anonymous):

hld on sec

OpenStudy (mertsj):

Discriminant is b^2-4ac

OpenStudy (mertsj):

\[ax^2+bx+c\] \[18x^2+11x+0\]

OpenStudy (mertsj):

a=18 b=11 c=0

OpenStudy (mertsj):

Calculate b^2-4ac

OpenStudy (usukidoll):

the discriminant formula is b^2-4ac find the values based on the the standard quadratic equation ax^2+bx+c let a = 18,b=11, c =0 and plug it into the formula

OpenStudy (anonymous):

47x

OpenStudy (usukidoll):

we should just have a number after calculating the discriminant

OpenStudy (anonymous):

47

OpenStudy (anonymous):

b² - 4ac = 11² - 4(18)(0)

OpenStudy (anonymous):

The part after the subtraction sign comes out to 0 so you're left with 11²

OpenStudy (usukidoll):

anything multiplied by 0 is 0 so we are left with 11^2-0 =11^2 what's 11 x 11 ?

OpenStudy (anonymous):

121

OpenStudy (usukidoll):

yes that's correct

OpenStudy (anonymous):

thanks very much